During a promotional stunt, Krusty the Clown stands on the roof of the Krusty Burger Restaurant and drops burgers down toward the hungry crowd below. One of the burgers is caught by Homer Simpson when its velocity is 12.50 m/s [D]. Calculate the burger’s displacement. 


1
Expert's answer
2021-01-11T11:40:51-0500

Let's first find the time that the burger takes to reach the ground from the kinematic equation:


"v_{y}=v_{0y}+gt,""t=\dfrac{v_y}{g}=\dfrac{12.5\ \dfrac{m}{s}}{9.8\ \dfrac{m}{s}}=1.275\ s."

Finally, we can find the vertical displacement of the burger (or the height of the Krusty Burger Restaurant) from another kinematic equation:


"y = \dfrac{1}{2}gt^2=\dfrac{1}{2}\cdot 9.8\ \dfrac{m}{s^2}\cdot(1.275\ s)^2=7.96\ m."

Answer:

"y = 7.96\ m."